KyledaGreat:

Solve the following exponential equation. Express your answer as both an exact expression and a decimal approximation rounded to two decimal places. 3e^4x = 63

KyledaGreat:

If you wish to enter log or ln, you must use the keypad. x = _____________ ≈ __________

Vocaloid:

is it 3e^(4x) = 6^3? or 63? or can you take a screencap?

KyledaGreat:

1 attachment
Vocaloid:

this one is one of the simpler ones divide both sides by 3 e^(4x) = 63/3 take the natural log of both sides to eliminate e, divide both sides by 4 to isolate x

KyledaGreat:

x = ln(21)/4

Vocaloid:

perfect for the decimal part, we can just chug it into a calculator and get x = 0.76113 which rounds to 0.76

KyledaGreat:

KyledaGreat:

Solve the following logarithmic equation. Express your answer as either an exact expression or a decimal approximation rounded to four decimal places. If there is no solution, indicate "No Solution (∅)." log_6(x−4)+log_6(x+1)=log_6(x+3) If you wish to enter log or ln, you must use the keypad. If there is more than one solution, separate your answers with a comma. Selecting a radio button will replace the entered answer value(s) with the radio button value. If the radio button is not selected, the entered answer is used. x = ____________________ No Solution (∅)

Vocaloid:

log rules - when you have two logs of the same base being added together, you can combine them, multiplying what's inside in other words: log_6(x−4)+log_6(x+1) ---> combine these two log base 6, multiply what's inside both logs ---> log_6( (x-4)*(x+1) ) setting both sides equal again log_6( (x-4)*(x+1) ) = log_6(x+3) notice how you have the same log base 6 on both sides, so you can raise both sides to the base of 6 and eliminate all the logs (x-4)(x+1) = (x+3) FOIL the left side, combine terms, either factor or use the quadratic formula to get your x-values

KyledaGreat:

\[x = 2 + {\sqrt{11}}\]

Vocaloid:

almost, for this quadratic equation we have 2 solutions (when you use the quadratic formula there's a +/- sign) x = 2 + sqrt(11) and x = 2 - sqrt(11)

KyledaGreat:

that's true

KyledaGreat:

it said wrong The answer you submitted, 2+√11,2−√11, is incorrect. It appears you have correctly used the quadratic formula to calculate possible solutions, but have made an error substituting these values back into log_6(x−4)+log_6(x+1)=log_6(x+3) to check them. You may wish to check your work. For a step-by-step guide to solving this problem, select Step By Step.

KyledaGreat:

i can still do it

Vocaloid:

ahh sorry, I forgot - logs can only take positive values in their input 2 - sqrt(11) doesn't work because it would make log(x+1) negative so it's only 2 + sqrt(11)

KyledaGreat:

it's okay

KyledaGreat:

correct

KyledaGreat:

another one is the same but different

KyledaGreat:

Solve the following logarithmic equation. Express your answer as either an exact expression or a decimal approximation rounded to four decimal places. If there is no solution, indicate "No Solution (∅)." log_4(x+1)−log_4(x−1)=2

KyledaGreat:

x = 17/15

KyledaGreat:

is that right

Vocaloid:

yup that's what I got too

KyledaGreat:

i knew it

KyledaGreat:

Solve the following exponential equation. Express your answer as both an exact expression and a decimal approximation rounded to two decimal places. 2^e2x + 10 = 97 x = _____________ ≈ __________ ln(97/2) / 2 -5 ≈ __________ -3.05921810

Vocaloid:

can you take a screengrab of the equation?

KyledaGreat:

ummm i don't think so , my phone acting stupid but i could try

Vocaloid:

can you try adding parentheses around the exponents? I just need to know what's in the exponent or not

KyledaGreat:

right , of course. excuse me one moment

KyledaGreat:

1 attachment
KyledaGreat:

i'm sorry if it's blurry , i had to use an ipad

KyledaGreat:

i tried to focus it

Vocaloid:

good just remember to round to 2 places -3.059 rounds up to -3.06

KyledaGreat:

Solve the following logarithmic equation. Express your answer as either an exact expression or a decimal approximation rounded to four decimal places. If there is no solution, indicate "No Solution (∅)." ln(x−4)=ln(5x)

Vocaloid:

they're both ln, so you can raise both sides to base e and eliminate the logs (x-4) = 5x, should be straightforward from there

KyledaGreat:

i think it's no solution

Vocaloid:

good, since x is negative ln(5x) becomes undefined, so that's correct

KyledaGreat:

Solve the following logarithmic equation. Express your answer as either an exact expression or a decimal approximation rounded to four decimal places. If there is no solution, indicate "No Solution (∅)." log_3(2x−5)+log_3(x)=log_3(5)

KyledaGreat:

\[x = \frac{ 5 + \sqrt{65}}{ 4}\]

Vocaloid:

good

KyledaGreat:

Solve the following exponential equation. Express your answer as both an exact expression and a decimal approximation rounded to two decimal places. Use e=2.71828182845905. e^2x+3 = 4^3x/7 x = _____________ ≈ __________

KyledaGreat:

-x = -21/14-ln(64) -2.13390412 are these right

Vocaloid:

yeah just make sure to round to -2.13

KyledaGreat:

Solve the following exponential equation. Express your answer as both an exact expression and a decimal approximation rounded to two decimal places. Use e=2.71828182845905. e^2x−2 = 132^4x+8 x = _____________ ≈ __________

KyledaGreat:

-x = -1+4 ln(132)/2ln(132) - 1 -2.34224681

Vocaloid:

I think you made a sign error in the left part should be -1 - 4ln(132) in the numerator other than that, good, make sure to round the decimal part -2.34

KyledaGreat:

okay you might be right

KyledaGreat:

no , i was right with the +

KyledaGreat:

Damon is saving up money for a down payment on a motorcycle. He currently has $5371, but knows he can get a loan at a lower interest rate if he can put down $6197. If he invests the $5371 in an account that earns 4.2% annually, compounded monthly, how long will it take Damon to accumulate the $6197? Round your answer to two decimal places, if necessary.

Vocaloid:

A = P(1 + r/n)^(nt) P is the principal (5371), r is the interest rate as a decimal (4.2%, so r = 0.042), n is the # of times compounded per year (monthly compounding, so n = 12), t is time (which you'll solve for). plug in and solve for t.

Vocaloid:

oh right, since he wants $6197, set A = 6197

KyledaGreat:

6197 years ?

Vocaloid:

no, t is time which is what you want to solve for. A is the amount he wants. set A = 6197 in the equation and solve for t.

KyledaGreat:

Oh sorry

KyledaGreat:

how could i set it up

KyledaGreat:

A = 6197 and 12 is the time right

Vocaloid:

the problem is asking how long the account will take to reach 6197, which means time is unknown, meaning you need to solve for t A = P(1 + r/n)^(nt) A = the final amount = 6197 P = principal = 5371 r = interest rate = 0.042 n = # of times compounded per year = 12 plug these in and solve for t

KyledaGreat:

5371(1 + 0.042/12)^(12t) ?

KyledaGreat:

is that right ?

Vocaloid:

yeah but you'd set that equal to the A-value, 6197 6197 = 5371(1 + 0.042/12)^(12t)

KyledaGreat:

oh okay

KyledaGreat:

t = 3.41193778

Vocaloid:

good just make sure to round to 3.41

KyledaGreat:

Tyron is saving up money for a down payment on a motorcycle. He currently has $3256, but knows he can get a loan at a lower interest rate if he can put down $3988. If he invests the $3256 in an account that earns 5.6% annually, compounded continuously, how long will it take Tyron to accumulate the $3988? Round your answer to two decimal places, if necessary.

Vocaloid:

continuous compounding is a bit different, there's a different formula A = Pe^(rt) A is the final amount (3988), P is the principal (3256), r is the interest rate (0.056), e is euler's constant, t is time, solve for t

KyledaGreat:

A = P(1 + r/n)^(nt) A = the final amount = 3988 P = principal = 3256 r = interest rate = 0.056

KyledaGreat:

3988 = 3256(1 + 0.056/12)^(12t)

Vocaloid:

this is a continuous compounding problem, so the formula is different

@vocaloid wrote:
continuous compounding is a bit different, there's a different formula A = Pe^(rt) A is the final amount (3988), P is the principal (3256), r is the interest rate (0.056), e is euler's constant, t is time, solve for t

KyledaGreat:

3988 = 3256e^(0.056t) is that right

Vocaloid:

yes

KyledaGreat:

t = 3.62125721

KyledaGreat:

is that right

Vocaloid:

yeah just be sure to round to 3.62

KyledaGreat:

In chemistry, the pH of a solution is a measure of the acidity or alkalinity of a solution. Water has a pH of 7 and, in general, acids have a pH less than 7 and alkaline solutions have a pH greater than 7. Find the pH of a solution with a hydronium ion concentration of 5.1×10^−6 moles/liter. Round your answer to two decimal places, if necessary. pH =

Vocaloid:

pH is -log(hydronium concentration), so it's just -log(5.1×10^−6), plug into a calculator and round appropriately